2. If 15.0 L Neon at 60 0C is heated at constant pressure, what temperature will it have if it occupies a volume of 4.50 L?

Given: V1= 15.0L V2 = 4.50 L T1 = 60 0C Convert 60 0C + 273 K = 333 K


1
Expert's answer
2022-05-16T17:02:16-0400

Given:

V1 = 15.0 L

T1 = 60°C + 273 = 333 K

P1 = P2 = constant

V2 = 4.50 L

T2 = unknown


Formula: V1/T1 = V2/T2


Solution:

Since the pressure and amount of gas remain unchanged, Charles's law can be used.

Charles's law can be expressed as: V1 / T1 = V2 / T2

V1T2 = V2T1

To find the temperture, solve the equation for T2:

T2 = V2T1 / V1

T2 = (4.50 L × 333 K) / (15.0 L) = 99.9 K = 100 K

T2 = 100 K or −173°C


Answer: at 100 K